OpenStudy (anonymous):

Which of following best describes the graph of the equations below? x – 2y = - 6 2x + y = 2 Select one: A. The lines have the same y-intercept. B. The lines have the same x-intercept. C. The lines are perpendicular. D. The lines are parallel.

OpenStudy (anonymous):

If you put the equations into a \[y = mx + b\] then you will find your answer

OpenStudy (anonymous):

y=2x+2 and y=-x/2-3 ? @hobbs

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

now if you look at the slopes you will find your answer

OpenStudy (anonymous):

wait how?

OpenStudy (anonymous):

We know that MX is our slope and we also see that the slopes are reciprocal (mean you flip the fractions and multiply by -1 ) of each other which usually means that they are perpendicular

OpenStudy (anonymous):

so its perpendicular?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you :)